OpenStudy (stephanos100):

Given that P = (-5, 5) and Q = (-13, 6), find the component form and magnitude of 2vector PQ. MEDALS!!!

OpenStudy (stephanos100):

is it <-16, 2>, square root of 130

OpenStudy (dumbcow):

component form is correct, magnitude is little off. \[|2PQ| = 2*\sqrt{(-8)^2 + 1^2} = 2 \sqrt{65}\]

OpenStudy (stephanos100):

There are two answers with that component form: one of my answers is <-16, 2>, square root of 130 and the other is <-16, 2>, square root of 260. Which would be correct? the first one, right?

OpenStudy (faiqraees):

square root of 260

OpenStudy (stephanos100):

whys that? thank you

OpenStudy (stephanos100):

so the correct answer is <-16, 2>, square root of 130

OpenStudy (faiqraees):

the options are like this component, magnitude?

OpenStudy (faiqraees):

because if that's the case then <-16, 2>, square root of 260 is correct

OpenStudy (stephanos100):

Okay thank you!

OpenStudy (faiqraees):

2√65 = √4 x √65 = √4x65 = √260